# Probabilistic Feature Augmentation for AIS-Based Multi-Path Long-Term Vessel Trajectory Forecasting

### v5

@@ Introduction
In Step 3 of our data processing pipeline, we discard messages within one kilometer of the nearest port as vessels are likely anchored in these areas. Additionally, we discard self-intersecting tracks and tracks that start and end at the same port, often observed in fishing vessels. After this filtering, we interpolate the remaining messages within each track to be 10 minutes apart based on their timestamp. Next, we remove tracks that are outliers from the dataset. Specifically, we identify outlier tracks within a grid cell where the number of tracks of a particular stratification pattern is less than or equal to five. That is the case of tracks of vessels patrolling around marine protected areas that may have been mislabeled during data collection and thus need to be removed. The resulting dataset consists of vessels grouped by their MMSI and divided by their type, with sets of tracks containing only coordinates. Finally, in the preprocessing stage, we reverse the tracks and add them to the dataset as new tracks to increase the sample size for the experiment. In step Step 4, we calculate the speed displacement of the vessel in knots and its acceleration in knots per hour based on the distance between the consecutive AIS messages. We do the same to compute the bearing of the vessel and the rate of the change of the bearing in gradian (see Methods). The resulting information is called the Standard Features of each track. Further, we split a track into two where the rate of change in the bearing between two consecutive points varies by more than ±$\pm 45^{g}$. The output of this process is divided into training, validation, and testing by vessel type and MMSI.
@@
Afterward, we split the dataset into three subsets: training, validation, and testing. The division is based on vessel type and MMSI, following the stratified sampling weights defined earlier. We use the training set to construct the probabilistic model. This model is capable of identifying common routes and predicting their probable destinations. The presence of multiple routes between frequently traveled source and destination ports highlights the need for creating route polygons to direct the forecasting models. In this sense, the set of Standard Features of each track is extended with the current grid cell, route polygon, and last grid cell centroids (Step 5). We refer to this extended feature set as Probabilistic Features. Subsequently to this step, in Step 6, we further extend the Probabilistic Features by calculating trigonometrical transformation on top of the vessels’ location and bearings resulting in the Trigonometrical Features. Either of the three feature sets can be used for training the deep learning model in Step T, but not multiple simultaneously. For example, in Figure 6, the Trigonometrical Features is the input of the deep learning model, whereas arrows from the other feature sets are greyed out.
@@
To generate smaller sequences of tracks for the deep learning model, we generate trajectories that are up to 15 hours long using a sliding window algorithm. This algorithm processes 3 new messages simultaneously, iterating over a track every 30 minutes. This aims to ensure that the input data is shaped into vectors of equal size, ranging from 1 to 3 hours, by padding any missing values. The output vectors always have a size of 12 hours, and no padding is applied to them. The deep learning model uses the first 3 hours of the output vector as input, while the remaining 12 hours are observed data used to evaluate the model’s prediction. We proceed to neural network engineering and training after generating the input features normalized between $[0,1]$. Regardless of the feature set we use, we expect the model to generate a pair of normalized coordinates in the range of $[0,1]$ as output, which, when denormalized, should resemble the anticipated output of the vessel trajectory. Such models are trained on the same training data used to build the probabilistic model, which is now segmented into trajectories. The training continues until no improvement is seen on the validation-reserved dataset slice, after which the models are evaluated on the test dataset. The training and validation are conducted simultaneously on tankers and cargo vessels. However, the final testing is performed separately so that we can account for the individual performance by vessel type. We also allow the last element in the training input vector to be present as the first element of the output vector (overlap of 10 minutes) to enhance the continuity of the sequential data.
@@
Our experiments involve testing different sets of features and network architectures. Our contributions are on feature engineering and neural network modeling, which together can make complex long-term path decisions. We conducted experiments with three distinct sets of features and displayed the results of other neural network models in an ablation fashion. We covered comparisons with the literature and state-of-art, including simple RNNs [38, 39, 40], Bidirectional RNNs [41, 42, 43], CNNs/RNNs AutoEncoders [44, 45, 46], and combinations of those with and without the transformer-based attention mechanism, all of which have less than 1.5M trainable parameters. We chose not to test models that exceeded the threshold due to their large size and long training time requirements. As AIS is a type of streaming data, it requires models that can train and produce fast outputs. Our proposal suggests using smarter features in combination with simpler, well-tailored models. Additionally, in the context of the Gulf of St. Lawrence, there is not enough data and information available to justify the use of larger and deeper models. These models may not necessarily provide better results but will require more resources and take longer sessions to train and refine.
@@
Differently from the literature, our model uses parallel Convolutional Neural Networks (CNNs) to extract local and global features from input data through dilated convolutions. The CNNs’ outputs are combined and processed by a Bidirectional Long Short-Term Memory (Bi-LSTM) with a Position-Aware Attention mechanism, efficiently encoding the temporal sequence, giving higher importance to later timesteps. The resulting latent space is then decoded through a second Bi-LSTM and multiple Dense layers, giving the final output as normalized latitude and longitude values. This end-to-end, customizable architecture achieves superior performance in predictive geospatial tasks, mainly when trained with the aid of probabilistic and trigonometrical features. To start this discussion, subsequently, we formalize our probabilistic model while using Table 1 to describe the symbols and notations used from hereinafter.
@@
Probabilistic Features and Probabilistic Model
A trajectory is a sequence of points representing the location of an object over time, and, in this paper, a trajectory $\tau$ is represented as an ordered sequence of $t$ points in time, such as $\tau=\left<x_{1},x_{2},\ldots,x_{t}\right>$, where a point $x_{t}\in\tau$ at time $t$ is also an ordered sequence of features from the AIS message such as $x_{t}=\left<\mathcal{L}_{on},\mathcal{L}_{at},v,\Delta_{v},\theta,\Delta_{\theta}\right>$; where $\mathcal{L}_{on}$ is the longitude, $\mathcal{L}_{at}$ the latitude, $v$ the speed in knots, $\theta$ the bearing in gradian, $\Delta_{v}$ the acceleration in knots per hour, and $\Delta_{\theta}$ the rate of change of the bearing in gradian. We refer to the set of values in $x_{t}$ of each AIS message as Standard Features. The speed, bearing, and deltas can be mathematically calculated using sequences of coordinates (see Methods for details).
@@
The similarity of a new trajectory $\tau_{n}$ and an existing historical track $\tau_{i}$ in a given grid cell $C_{i}$ is computed using Euclidean distance of their motion statistics $S^{C_{i}}_{\tau_{n}}$ and $S^{C_{i}}_{\tau_{i}}$, respectively, in $C_{i}$ using equation:
@@
$\hfill\mathbb{P}(R_{k}|C_{i},\tau_{n})=\frac{\sum_{\tau_{i}\in T}\mathbf{I}\Big{(}x_{t}\in\tau_{i}\wedge\exists x_{t}\in R_{k}\,\wedge\,E(S_{\tau_{n}},S_{\tau_{i}})\leq\delta\,\wedge\,\exists x_{t}\in C_{i}\Big{)}}{\sum_{\tau_{i}\in T}\mathbf{I}\Big{(}x_{t}\in\tau_{i}\,\wedge\,E(S_{\tau_{n}},S_{\tau_{i}})\leq\delta\,\wedge\,\exists x_{t}\in C_{i}\Big{)}};\hfill$ (2)
@@
$\hfill\mathbb{P}(D_{j}|C_{i},R_{k},\tau_{n})=\frac{\sum_{\tau_{i}\in T}\mathbf{I}\Big{(}x_{t}\in\tau_{i}\,\wedge\,\exists x_{t}\in D_{j}\,\wedge\,E(S_{\tau_{n}},S_{\tau_{i}})\leq\delta\,\wedge\,\exists x_{t}\in C_{i}\,\wedge\,\exists x_{t}\in R_{k}\Big{)}}{\sum_{\tau_{i}\in T}\mathbf{I}\Big{(}x_{t}\in\tau_{i}\,\wedge\,E(S_{\tau_{n}},S_{\tau_{i}})\leq\delta\,\wedge\,\exists x_{t}\in C_{i}\,\wedge\,\exists x_{t}\in R_{k}\Big{)}};\hfill$ (3)
@@
$\hfill\mathbb{P}(D_{j}|C_{i},\tau_{n})=\frac{\sum_{\tau_{i}\in T}\mathbf{I}\Big{(}x_{t}\in\tau_{i}\,\wedge\,\exists x_{t}\in D_{j}\,\wedge\,E(S_{\tau_{n}},S_{\tau_{i}})\leq\delta\,\wedge\,\exists x_{t}\in C_{i}\Big{)}}{\sum_{\tau_{i}\in T}\mathbf{I}\Big{(}x_{t}\in\tau_{i}\,\wedge\,E(S_{\tau_{n}},S_{\tau_{i}})\leq\delta\,\wedge\,\exists x_{t}\in C_{i}\Big{)}};\hfill$ (4)
@@
$\gamma^{\leavevmode\nobreak\ d}$, $\beta^{d}$ Scaling and shifting parameters for each filter $d$ in batch normalization
@@
For speed, applying the logarithmic transformation can be effective, especially when dealing with values spanning across multiple orders of magnitude. This can ensure that larger-scale patterns do not overly influence the model, allowing local trends to play a part in the decision-making process. This can be achieved with the equation $v^{\prime}=\log(1+|v|)$, where $v^{\prime}$ is the transformed and $v$ is the original speed. Through these sets of transformations, the earlier point $x$ at time $t$ from a trajectory $\tau$ become $x_{t}=\Big{<}\mathcal{L}_{on},\mathcal{L}_{at},\alpha,\beta,\gamma,v,\Delta_{v},\theta,\Delta_{\theta},v^{\prime},\Delta_{v^{\prime}},\cos_{\theta},\sin_{\theta},\cos_{\Delta_{\theta}},\sin_{\Delta_{\theta}},\mathcal{L}_{on}^{R},\mathcal{L}_{at}^{R},\alpha^{R},\beta^{R},\gamma^{R},\mathcal{L}_{on}^{N},\mathcal{L}_{at}^{N},\alpha^{N},\beta^{N},\gamma^{N},$ $\mathcal{L}_{on}^{D},\mathcal{L}_{at}^{D},\alpha^{D},\beta^{D},\gamma^{D}\Big{>}$. This final representation includes Standard, Probabilistic, and Trigonometrical Features. It provides an enriched input dataset for the deep-learning model tailored for trajectory forecasting, facilitating its understanding of complex spatial and temporal patterns of the trajectories. With all feature sets properly defined, we move forward with the formalization of the deep-learning model (see Figure 9), which is based on Table 2 for describing the symbols and notations used from hereinafter.
Deep Learning Model
The deep learning model comprises two main parts; spatial and temporal feature learning, as depicted in Figure 9. Our customized architecture uses parallel convolutional neural networks with recurrent neural networks and focal attention for position-aware temporal learning. During the model formulation, we use lower-case bold letters to represent vectors, such as $\mathbf{x}$, upper-case bold letters to represent matrices, such as $\mathbf{W}$, and non-bold lower-case values, such as $b$, to indicate a scalar value or an integer index of matrices and vectors. The equations are defined with indices to help understand how the data changes as it flows through the architecture and to indicate the meaning of each process accomplished by the network module. We use $\mathbf{x}$ to describe the input tensor, which is propagated and updated at each process step. The weights and biases, $\mathbf{W}$ and $\mathbf{b}$, are independent of each other but are optimized together during the network training.
@@
$\hfill\mathbf{x}_{b,i,m}=\mathbf{max}\left(\frac{\mathbf{x}_{b,p,m}-\mu_{b}^{d}}{\sqrt{\big{(}\sigma_{b}^{d}\big{)}^{2}+\epsilon}}\cdot\gamma^{\leavevmode\nobreak\ d}+\beta^{d}\right)\quad\forall p\in[i,i+h-1]\hfill$ (9)
For each filter $m$ and batch $b$, a window of size $h$ is defined on the feature map starting at position $i$, with the condition that $i+h-1\leq\widetilde{w}$. The filter-wise mean $\mu_{b}^{d}$ and variance $\big{(}\sigma_{b}^{d}\big{)}^{2}$ are calculated for batch $b$, along with specific scaling $\gamma^{\leavevmode\nobreak\ d}$ and shifting $\beta^{d}$ parameters. Next, the max pooling operation is performed over the window, extracting the maximum value of $p$. The resulting value is assigned to $\mathbf{x}_{b,i,d}$, which becomes the feature map for the given window of filter index $d$. Due to working with parallel convolutions that involve different kernel sizes, it is necessary to use a combination of techniques to ensure consistent tensor dimensions. We use batch normalization and max-pooling to scale down the tensor dimensions into shared tensor dimensions with scaled values, which can be merged later. This approach guarantees that the branches can be combined seamlessly without stability issues during the training.
As mentioned, single-dimensional CNNs can only learn spatial features within their kernel size. We employ parallel convolutions in residual blocks to capture both short-term and long-term spatial features simultaneously. This is achieved through dilated convolutional layers, which expand the kernel size by a factor of $\vartheta$ and enable it to capture extended spatial features from the input data. Both convolutional branches in a block are trained together, but they work independently on copies of the input data so that they are not affected by each other. To perform dilated convolution, we modify Equation 8 as follows:
@@
where $j$ stands for the filter index, and $\vartheta\times(j-1)$ is the expansion of the kernel due to the dilation rate $\vartheta$. The resulting output of the branches, given by Equation 9, is then combined into a single tensor as such:
@@
where $T$ and $D$ denote the results obtained from the traditional and dilated convolution operations.
The series of mathematical operations that end in Equation 11, as shown in Figure 9, is repeated three times, each time with different parameters. At this point, $\widetilde{w}$ and $m$ have dimensions according to the hyperparameters of the convolution block (see Methods). This stack allows us to extract precise long and short-term spatial features from the input data at each process step. The parallel CNNs and the sequential blocks work simultaneously to produce a refined output of spatial features that is then sent to the next block for further processing, minimizing further the overall error for the forecasting task.
The subsequent part is where the network learns the temporal features from the input data. It starts with a dropout layer to prevent overfitting from the spatial feature data. The dropout layer comprises a binary mask of shape $\mathbb{R}^{b\times\widetilde{w}\times m}$, where each element has a value of 1 with probability $p$ and a value of 0 with probability $1-p$, which is element-wise multiplied with the input data to generate the output. Such output is forwarded to a Bidirectional Long-Short Term Memory (Bi-LSTM) network, which takes $\mathbf{x}\in\mathbb{R}^{b\times\widetilde{w}\times m}$ as input and produces the hidden state $\mathbf{h}_{b,i,r}\in\mathbb{R}^{b\times\widetilde{w}\times r}$ for each time step $i\leq\widetilde{w}$, where $r$ is the dimension of the LSTM inner weights. The Bi-LSTM equations can be formulated as follows:
@@
where $\sigma$ is the sigmoid function, $\mathbf{tanh}$ is the hyperbolic tangent function, $[\mathbf{h}_{b,(i-1),r}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \mathbf{x}_{b,i,m}]$ is the concatenation of the previous hidden state with the $i$-th element of the input. $\mathbf{i}$, $\mathbf{f}$, $\widetilde{\mathbf{C}}$, $\mathbf{o}$, $\mathbf{C}$, and $\mathbf{h}$ represent the input gate, forget gate, candidate cell state, output gate, cell state, and hidden state, respectively. $\mathbf{x}_{b,i,r}$ is the output of the process, where $\mathbf{W}_{d}$ and $\mathbf{b}_{d}$ is the weight and bias for the dense layer after the Bi-LSTM and $\phi$ is the ReLU activation. Because Bi-LSTMs compute the hidden states in the forward $\mathbf{h}_{b,i,r}^{\rightarrow}$ and backward $\mathbf{h}_{b,i,r}^{\leftarrow}$ direction, the double output is merged before going through the fully-connected layer.
@@
Results
The evaluation of results begins with the probabilistic features, which constitute the foundation of our proposal and the main catalyst for improved decision-making using neural networks in the task of multi-path long-term vessel trajectory forecasting. The probability model is trained on the same dataset that will be later used to train the deep-learning model. The crucial difference is that the deep-learning model’s success hinges on the probabilistic model’s successful implementation. The probabilistic model is responsible for feature engineering, while the deep-learning model focuses on trajectory forecasting based on the vessel position and its relevant probabilistic features. This means that when the deep-learning model has access to data regarding the vessel’s probable route and destination, it can significantly decrease the uncertainty in trajectory forecasting, leading to better performance for the forecasting pipeline.

### v6

@@ 2 Problem Statement & Methodology
In Step 3 of our data processing pipeline, we discard messages within one kilometer of the nearest port as vessels are likely anchored in these areas. Additionally, we discard self-intersecting tracks as well as tracks that start and end at the same port, often observed in fishing vessels. After this filtering, we interpolate points of each track with uniform interval of 10 minutes between two consecutive points. Next, we remove tracks that are outliers from the dataset. Specifically, we identify outlier tracks within a grid cell where the number of tracks of a particular stratification pattern is less than or equal to five. That is the case of tracks of vessels patrolling around marine protected areas that may have been mislabeled during data collection and thus need to be removed. The resulting dataset consists of vessels grouped by their MMSI and divided by their type, with sets of tracks containing only coordinates. Finally, in the preprocessing stage, we reverse the tracks and add them to the dataset as new tracks to increase the sample size for the experiment. In step Step 4, we calculate the speed displacement of the vessel in knots and its acceleration in knots per hour based on the distance between the consecutive AIS messages. We do the same to compute the bearing of the vessel and the rate of the change of the bearing in gradian (see Appendix). The resulting information is called the Standard Features (i.e., longitude, latitude, speed, bearing, change in speed, change in bearing) of each track.
@@
Afterward, we split the dataset into three subsets: training, validation, and testing. The division is based on vessel type and MMSI, following the stratified sampling weights defined earlier. We use the training set to construct the probabilistic model. This model is capable of identifying common routes and predicting their probable destinations. The presence of multiple routes between frequently traveled source and destination ports highlights the need to create route polygons to direct the forecasting models. In this sense, the set of Standard Features of each track is extended with the current grid cell, route polygon, and last grid cell centroids (Step 5). We refer to this extended feature set as Probabilistic Features. Subsequently to this step, in Step 6, we further extend the Probabilistic Features by calculating trigonometrical transformation on top of the vessels’ location and bearings resulting in the Trigonometrical Features. Either of the three feature sets can be used for training the deep learning model in Step T, but not multiple simultaneously. For example, in Figure 6, the Trigonometrical Features is the input of the deep learning model, whereas arrows from the other feature sets are greyed out.
@@
The target of our approach is to forecast long term trajectory up to 12 hours by feeding 3 hours trajectory to the model. The tracks of the vessels are usually in days so smaller sequences (sub-tracks) are needed. To generate smaller sequences of tracks for the deep learning model, we generate trajectories that are up to 15 hours long using a sliding window algorithm. This algorithm processes 3 new messages simultaneously, iterating over a track every 30 minutes. This aims to ensure that the input data is shaped into vectors of equal size, ranging from 1 to 3 hours, by padding any missing values. The output vectors always have a size of 12 hours, and no padding is applied to them. The deep learning model uses the first 3 hours of the output vector as input, while the remaining 12 hours are observed data used to evaluate the model’s prediction. We proceed to neural network engineering and training after generating the input features normalized between $[0,1]$. Regardless of the feature set we use, we expect the model to generate a pair of normalized coordinates in the range of $[0,1]$ as output, which, when denormalized, should resemble the anticipated output of the vessel trajectory. It is important to mention that probabilistic model is learned with tracks in training dataset whereas deep learning model is trainined over the trajectories (sub-tracks or subsequences). The training of deep learning model continues until no improvement is seen on the validation-reserved dataset slice, after which the models are evaluated on the test dataset. The training and validation are conducted simultaneously on tankers and cargo vessels. However, the final testing is performed separately so that we can account for the individual performance by vessel type. We also allow the last element in the training input vector to be present as the first element of the output vector (overlap of 10 minutes).
@@
Our experiments involve testing different sets of features and network architectures. Our contributions are on feature engineering and neural network modeling, which together can make complex long-term path decisions. We conducted experiments with three distinct sets of features and displayed the results of other neural network models in an ablation fashion. We covered comparisons with the literature and state-of-art, including simple RNNs [38, 39, 40], Bidirectional RNNs [41, 42, 43], CNNs/RNNs AutoEncoders [44, 45, 46], and combinations of those with and without the transformer-based attention mechanism, all of which have less than 1.5M trainable parameters. We chose not to test models that exceeded the threshold due to their large size and long training time requirements. As AIS is a type of streaming data, it requires models that can train and produce fast outputs. Our proposal suggests using smarter features in combination with simpler, well-tailored models. Additionally, in the Gulf of St. Lawrence context, there is not enough data and information available to justify using larger and deeper models. These models may not necessarily provide better results but will require more resources and take longer training sessions.
@@
Differently from the literature, our model uses parallel Convolutional Neural Networks (CNNs) to extract local and global features from input data through dilated convolutions. The CNNs’ outputs are combined and processed by a Bidirectional Long Short-Term Memory (Bi-LSTM) with a Position-Aware Attention mechanism, efficiently encoding the temporal sequence, giving higher importance to later timesteps. The resulting latent space is then decoded through a second Bi-LSTM and multiple Dense layers, giving the final output as normalized latitude and longitude values. This end-to-end, customizable architecture achieves superior performance in predictive geospatial tasks, mainly when trained with probabilistic and trigonometrical features. To start this discussion, subsequently, we formalize our probabilistic model while using Table 1 to describe the notations used from hereinafter.
@@
3 Probabilistic Feature Model
A trajectory is a sequence of points representing the location of an object over time, and, in this paper, a trajectory $\tau$ is represented as an ordered sequence of $t$ points in time, such as $\tau=\left<x_{1},x_{2},\ldots,x_{t}\right>$, where a point $x_{t}\in\tau$ at time $t$ is also an ordered sequence of features from the AIS message such as $x_{t}=\left<\mathcal{L}_{on},\mathcal{L}_{at},v,\Delta_{v},\theta,\Delta_{\theta}\right>$; where $\mathcal{L}_{on}$ is the longitude, $\mathcal{L}_{at}$ the latitude, $v$ the speed in knots, $\theta$ the bearing in gradian, $\Delta_{v}$ the acceleration in knots per hour, and $\Delta_{\theta}$ the rate of change of the bearing in gradian. We refer to the set of values in $x_{t}$ of each AIS message as Standard Features. The speed, bearing, and deltas can be mathematically calculated using sequences of coordinates (see Annex for details).
@@ 3 Probabilistic Feature Model
The similarity of a new trajectory $\tau_{n}$ and a historical track $\tau_{i}$ in a given grid cell $C_{i}$ is computed using Euclidean distance of their motion statistics $S^{C_{i}}_{\tau_{n}}$ and $S^{C_{i}}_{\tau_{i}}$ in $C_{i}$ as follows:
@@
$\hfill\mathbb{P}(R_{k}|C_{i},\tau_{n})=\frac{\sum_{\tau_{i}\in T}\mathbf{I}\Big{(}x_{t}\in\tau_{i}\wedge\exists x_{t}\in R_{k}\,\wedge\,E(S_{\tau_{n}},S_{\tau_{i}})\leq\delta\,\wedge\,\exists x_{t}\in C_{i}\Big{)}}{\sum_{\tau_{i}\in T}\mathbf{I}\Big{(}x_{t}\in\tau_{i}\,\wedge\,E(S_{\tau_{n}},S_{\tau_{i}})\leq\delta\,\wedge\,\exists x_{t}\in C_{i}\Big{)}}\hfill$ (2)
@@
$\mathbb{P}(D_{j}|C_{i},R_{k},\tau_{n})=\frac{\sum_{\tau_{i}\in T}\mathbf{I}\Big{(}x_{t}\in\tau_{i}\,\wedge\,\exists x_{t}\in D_{j}\,\wedge\,E(S_{\tau_{n}},S_{\tau_{i}})\leq\delta\,\wedge\,\exists x_{t}\in C_{i}\,\wedge\,\exists x_{t}\in R_{k}\Big{)}}{\sum_{\tau_{i}\in T}\mathbf{I}\Big{(}x_{t}\in\tau_{i}\,\wedge\,E(S_{\tau_{n}},S_{\tau_{i}})\leq\delta\,\wedge\,\exists x_{t}\in C_{i}\,\wedge\,\exists x_{t}\in R_{k}\Big{)}}\hfill$ (3)
@@
$\hfill\mathbb{P}(D_{j}|C_{i},\tau_{n})=\frac{\sum_{\tau_{i}\in T}\mathbf{I}\Big{(}x_{t}\in\tau_{i}\,\wedge\,\exists x_{t}\in D_{j}\,\wedge\,E(S_{\tau_{n}},S_{\tau_{i}})\leq\delta\,\wedge\,\exists x_{t}\in C_{i}\Big{)}}{\sum_{\tau_{i}\in T}\mathbf{I}\Big{(}x_{t}\in\tau_{i}\,\wedge\,E(S_{\tau_{n}},S_{\tau_{i}})\leq\delta\,\wedge\,\exists x_{t}\in C_{i}\Big{)}}\hfill$ (4)
@@
$\gamma^{~{}d}$, $\beta^{d}$ Scaling and shifting parameters for each filter $d$ in batch normalization
@@
For speed, applying the logarithmic transformation can be effective, especially when dealing with values spanning across multiple orders of magnitude. This can ensure that larger-scale patterns do not overly influence the model, allowing local trends to play a part in the decision-making process. This can be achieved with the equation $v^{\prime}=\log(1+|v|)$, where $v^{\prime}$ is the transformed and $v$ is the original speed. Through these sets of transformations, the earlier point $x$ at time $t$ from a trajectory $\tau$ become $x_{t}=\Big{<}\mathcal{L}_{on},\mathcal{L}_{at},\alpha,\beta,\gamma,v,\Delta_{v},\theta,\Delta_{\theta},v^{\prime},\Delta_{v^{\prime}},\\ \cos_{\theta},\sin_{\theta},\cos_{\Delta_{\theta}},\sin_{\Delta_{\theta}},\mathcal{L}_{on}^{R},\mathcal{L}_{at}^{R},\alpha^{R},\beta^{R},\gamma^{R},\mathcal{L}_{on}^{N},\mathcal{L}_{at}^{N},\alpha^{N},\beta^{N},\gamma^{N},$ $\mathcal{L}_{on}^{D},\mathcal{L}_{at}^{D},\alpha^{D},\beta^{D},\gamma^{D}\Big{>}$. This final representation includes Standard, Probabilistic, and Trigonometrical Features. It provides an enriched input dataset for the deep-learning model tailored for trajectory forecasting, facilitating its understanding of complex spatial and temporal patterns of the trajectories. With all feature sets properly defined, we move forward with the formalization of the deep-learning model (see Figure 9), which is based on Table 2 for describing the symbols and notations used from hereinafter.
4 Deep Learning Model
The deep learning model comprises two main parts; spatial and temporal feature learning, as depicted in Figure 9. Our customized architecture uses parallel CNN with recurrent neural networks and focal attention for position-aware temporal learning. During the model formulation, we use lower-case bold letters to represent vectors, such as $\mathbf{x}$, upper-case bold letters to represent matrices, such as $\mathbf{W}$, and non-bold lower-case values, such as $b$, to indicate a scalar value or an integer index of matrices and vectors. The equations are defined with indices to help understand how the data changes as it flows through the architecture and to indicate the meaning of each process accomplished by the network module. We use $\mathbf{x}$ to describe the input tensor, which is propagated and updated at each process step. The weights and biases, $\mathbf{W}$ and $\mathbf{b}$, are independent of each other but are optimized together during the network training.
@@ 4 Deep Learning Model
$\hfill\mathbf{x}_{b,i,m}=\mathbf{max}\left(\frac{\mathbf{x}_{b,p,m}-\mu_{b}^{d}}{\sqrt{\big{(}\sigma_{b}^{d}\big{)}^{2}+\epsilon}}\cdot\gamma^{~{}d}+\beta^{d}\right)\quad\forall p\in[i,i+h-1]\hfill$ (9)
For each filter $m$ and batch $b$, a window of size $h$ is defined on the feature map starting at position $i$, with the condition that $i+h-1\leq\widetilde{w}$. The filter-wise mean $\mu_{b}^{d}$ and variance $\big{(}\sigma_{b}^{d}\big{)}^{2}$ are calculated for batch $b$, along with specific scaling $\gamma^{~{}d}$ and shifting $\beta^{d}$ parameters. Next, the max pooling operation is performed over the window, extracting the maximum value of $p$. The resulting value is assigned to $\mathbf{x}_{b,i,d}$, which becomes the feature map for the given window of filter index $d$. Due to working with parallel convolutions that involve different kernel sizes, it is necessary to use a combination of techniques to ensure consistent tensor dimensions. We use batch normalization and max-pooling to scale down the tensor dimensions into shared tensor dimensions with scaled values, which can be merged later. This approach guarantees that the branches can be combined seamlessly without stability issues during the training.
As mentioned, single-dimensional CNNs can only learn spatial features within their kernel size. We employ parallel convolutions in residual blocks to simultaneously capture short-term and long-term spatial features. This is achieved through dilated convolutional layers, which expand the kernel size by a factor of $\vartheta$ and enable it to capture extended spatial features from the input data. Both convolutional branches in a block are trained together, but they work independently on copies of the input data so that they are not affected by each other. To perform dilated convolution, we modify Equation 8 as follows:
@@
where $j$ is the filter index, and $\vartheta\times(j-1)$ is the expansion of the kernel due to the dilation rate $\vartheta$. The output from Equation 9 is then combined into a single tensor as such:
@@
where $T$ and $D$ denote the results obtained from the traditional and dilated convolution operations. Notice that $\widetilde{w}$ and $m$ are updated each time they go through the CNN blocks.
The series of mathematical operations that end in Equation 11, as shown in Figure 9, is repeated three times, each time with different parameters. At this point, $\widetilde{w}$ and $m$ have dimensions according to the hyperparameters of the convolution block (see Annex). This stack allows us to extract precise long and short-term spatial features from the input data at each process step. The parallel CNNs and the sequential blocks work simultaneously to produce a refined output of spatial features that is then sent to the next block for further processing, minimizing further the overall error for the forecasting task.
The subsequent part is where the network learns the temporal features from the input data. It starts with a dropout layer to prevent overfitting from the spatial feature data. The dropout layer comprises a binary mask of shape $\mathbb{R}^{b\times\widetilde{w}\times m}$, where each element has a value of 1 with probability $p$ and a value of 0 with probability $1-p$, which is element-wise multiplied with the input data to generate the output. Such output is forwarded to a Bidirectional Long-Short Term Memory (Bi-LSTM) network, which takes $\mathbf{x}\in\mathbb{R}^{b\times\widetilde{w}\times m}$ as input and produces the hidden state $\mathbf{h}_{b,i,r}\in\mathbb{R}^{b\times\widetilde{w}\times r}$ for each time step $i\leq\widetilde{w}$, where $r$ is the dimension of the LSTM inner weights. The Bi-LSTM equations can be formulated as:
@@
where $\sigma$ is the sigmoid function, $\mathbf{tanh}$ is the hyperbolic tangent function, $[\mathbf{h}_{b,(i-1),r}~{},~{}\mathbf{x}_{b,i,m}]$ is the concatenation of the previous hidden state with the $i$-th element of the input. $\mathbf{i}$, $\mathbf{f}$, $\widetilde{\mathbf{C}}$, $\mathbf{o}$, $\mathbf{C}$, and $\mathbf{h}$ represent the input gate, forget gate, candidate cell state, output gate, cell state, and hidden state, respectively. $\mathbf{x}_{b,i,r}$ is the output of the process, where $\mathbf{W}_{d}$ and $\mathbf{b}_{d}$ is the weight and bias for the dense layer after the Bi-LSTM and $\phi$ is the ReLU activation. Because Bi-LSTMs compute the hidden states in the forward $\mathbf{h}_{b,i,r}^{\rightarrow}$ and backward $\mathbf{h}_{b,i,r}^{\leftarrow}$ direction, the double output is merged before going through the fully-connected layer.
@@
5 Results
The evaluation of results begins with the probabilistic features, which constitute the foundation of our proposal and the main catalyst for improved decision-making using neural networks in the task of multi-path long-term vessel trajectory forecasting. The probability model is trained on the same dataset that is later used to train the deep-learning model. The crucial difference is that the deep-learning model’s success hinges on the probabilistic model’s successful implementation. The probabilistic model is responsible for adding route and destination information in standard feature set, while the deep-learning model focuses on trajectory forecasting based on the vessel position and its relevant probabilistic features. This means that when the deep-learning model has access to data regarding the vessel’s probable route and destination, it can significantly decrease the uncertainty in trajectory forecasting, leading to better performance for the forecasting pipeline.